how do you solve -3/u+1=-8+7/u+2
Step 1: Group like terms together
-3/u-7/u = -8+2-1
Step 2: find LCF of the demoninators of the equation at the left of the equal sign. Here you will realise that U is the LCF and so factorised as
-3-7/U = -7
Step 3: Solve for U value
-10/U = -7
Mutliply both sides by U
(U)-10/U = -7(U)
The two Us at left will cancel themselves and you will have:
-10 = -7U
Divide both sides by -7:
-10/-7 = -7U/-7
At the left the negatives will cancel themselves and at the right -7 as a demoninator will cancel the numerator.
Therefore, the final answer will be:
10/7 = U or U = 1 whole number, 3 over 7 (Mixed fraction)



Answer by freekie11(2)

Step 1: Group like terms together
-3/u-7/u = -8+2-1
Step 2: find LCF of the demoninators of the equation at the left of the equal sign. Here you will realise that U is the LCF and so factorised as
-3-7/U = -7
Step 3: Solve for U value
-10/U = -7
Mutliply both sides by U
(U)-10/U = -7(U)
The two Us at left will cancel themselves and you will have:
-10 = -7U
Divide both sides by -7:
-10/-7 = -7U/-7
At the left the negatives will cancel themselves and at the right -7 as a demoninator will cancel the numerator.
Therefore, the final answer will be:
10/7 = U or U = 1 whole number, 3 over 7 (Mixed fraction)